OpenStudy (anonymous):

solve quadratic inequalities: x2-2x is greater than or equals to 0

OpenStudy (anonymous):

are you missing a ^ sign , i.e. is it x squared - 2*x

OpenStudy (anonymous):

Oh its x^2-2x>=0 right?

OpenStudy (anonymous):

Can u convert x^2-2x into completed square form

OpenStudy (anonymous):

its x^2-2x>=0

OpenStudy (anonymous):

find the zeros. it is negative between them and positive outside them because it is a parabola that opens up

OpenStudy (anonymous):

x^2-2x>=0 x^2-2*x*1+1^2>=1^2 (x-1)^2>=1

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

Now, HINT: IF A^2>=(+-b)^2 then A<=-b AND A>=b

OpenStudy (anonymous):

lets try this \[x^2-2x\geq0\] \[x(x-2)\geq 0\] zeros are at \(x=0\) and at \(x=2\) it will be negative between the zeros, i.e. on the interval \((0,2)\) and positive outside the zeros, on the intervals \[(-\infty,0)\cup (2,\infty)\]

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

\(y=x^2-2x\) is a parabola that open up 

OpenStudy (anonymous):

it is negative (below the \(x\) axis) between the zeros positive (above the \(x\) axis) outside the zeros

OpenStudy (anonymous):

@satellite73 thank you

OpenStudy (anonymous):

This may also help: x^2-2x>=0 x^2-2*x*1+1^2>=1^2 (x-1)^2>=1 (x-1)^2>=(+-1)^2 Now, x-1<=-1 AND x-1>=1 x<=0 AND x>=2

OpenStudy (anonymous):

not to be picky, but there is no such thing as \(x\leq 0\) AND \(x\geq 2\) it should be \[ x\leq 0 \text{ or } x\geq 2\]